We're currently investigating an issue with a switch that has affected access to a small number of cloud servers.
We're power cycling the affected switch at the moment and expect it to come back online. ETA currently 15 minutes.
The power cycle of the affected switch is now complete and traffic is passing normally to the affected servers. We are continuing to investigate the issue.